# DRAT Proofs of Unsatisfiability for SAT Modulo Monotonic Theories

### 19 Jan 2024 v1

@@ 1 Introduction
  • DRAT is a desirable proof format. (Here, we include related formats like DRUP [27], GRIT [19], and LRAT [18]. DRAT is explained in Sec. 2.1.) For an independent assurance of correctness, the proof checker is the critical, trusted component, and hence must be as trustworthy as possible. For (propositional) SAT, the community has coalesced around the DRAT proof format [38], for which there exist independent, efficient proof checkers [38], mechanically verified proof checkers [39], and even combinations that are fast as well as mechanically proven [18]. Since 2014, in order to participate in the main track of the annual SAT Competition, SAT solvers are required to produce checkable DRAT certificate for unsatisfiable instances.
@@
We implemented our method in the MonoSAT SMMT solver [10]. For evaluation, we use two sets of benchmarks derived from practical, industrial problems: multilayer escape routing [11], and cloud network reachability [2]. We will make both sets of benchmarks available publicly. Our results show minimal runtime overhead on the solver (geometric mean slowdown 7.4%, worst-case 28.8% in our experiments), and we generate and check proofs for many problem instances that are otherwise intractable.
@@ 2.1 Propositional SAT and DRAT
DRAT [38] is a much more compact and efficient system for proving unsatisfiability. It is based on reverse unit propagation (RUP), which we explain here.111 RUP is all we use in this paper. RAT is a superset of RUP, by essentially doing one step of resolution as a “lookahead” before checking RUP of the resolvents. The “D” in DRAT stands for “deletion”, meaning the proof format also records clause deletions. A unit clause is a clause containing one literal. If $L$ is the set of literals appearing in the unit clauses of a formula $F$, the unit clause rule computes $F|_{L}$, and the repeated application of the unit clause rule until a fixpoint is called unit propagation (aka Boolean constraint propagation). Given a clause $C$, its negation $\overline{C}$ is a set of unit clauses, and we denote by $F\vdash_{1}C$ if $F\wedge\overline{C}$ derives a conflict through unit propagation. Notice that $F\vdash_{1}C$ implies $F\models C$, but is computationally easy to check. The key insight [24] behind RUP is that modern CDCL SAT solvers make progress by deriving learned clauses, whose redundancy is, by construction, checkable via unit propagation. Proof generation, therefore, is essentially just logging the sequence of learned clauses leading to $\bot$, and proof checking is efficiently checking $\vdash_{1}$ of the relevant learned clauses.
@@ 2.2 SAT Modular Monotonic Theories (SMMT)
Given a monotonic predicate $p$ over input $A$, we will use boldface $\mathbf{p}$ as the predicate atom for $p$, i.e., the predicate atom would be a Boolean variable in the CNF encoding of the theory, indicating whether $p(A)$ is true or not. The theory of $\mathbf{p}$ is the set of valid implications in the form of $\mathcal{M}_{A}\Rightarrow\mathbf{p}$ where $\mathcal{M}_{A}$ is a partial assignment over $A$.
@@ 4.2 Monotonic Definitions
The proof of Theorem 4.2 is in Appendix 0.A.1. The correctness relies on the fact that the predicate $p$ is indeed monotonic, and that our propositional definitions need only be one-sided. If the monotonic definition is already in Horn theory, it can be used directly verify theory lemmas via RUP; otherwise, we proceed to Horn approximation, described next.
@@ 4.3 Instantiation-Based, Proof-Specific Horn Definition
Let $p$ be a monotonic predicate over Boolean arguments $A$. Suppose $\Sigma_{\mathbf{p}}$ and $\Sigma_{\bar{\mathbf{p}}}$ are the positive and negative propositional definitions, respectively. For every assignment $\mathcal{M}$ to the variables in $A$:
@@
The proof of Lemma. 1 is presented in the Appendix. 0.A.2. The duality of the positive ($\Sigma_{\mathbf{p}}$) and negative ($\Sigma_{\bar{\mathbf{p}}}$) definitions allows us to over-approximate positive (negative) definitions by instantiating the negative (positive) definitions.
@@
Let a predicate $p$ over input $A$ and a positive definition $\Sigma_{\mathbf{p}}$ be given. For any partial assignment $M^{\prime}$ over $var(\Sigma_{\mathbf{p}})\setminus(var(p)\cup A)$, $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\Rightarrow\overline{\mathbf{p}}$ is an over-approximation of $\Sigma_{\overline{\mathbf{p}}}$.    222 Note that $\Sigma_{\mathbf{p}}|_{M^{\prime}}$ is encoded in CNF, so to compactly (i.e., linear-size) encode $\Sigma_{\mathbf{p}}|_{M^{\prime}}\Rightarrow\overline{\mathbf{p}}$ in CNF, we introduce a new literal $l_{i}$ for each clause $C_{i}\in\Sigma_{\mathbf{p}}|_{M^{\prime}}$, create clauses $\overline{c_{ij}}\vee l_{i}$ for each literal $c_{ij}\in C_{i}$, and add clause $\overline{l_{1}}\vee\overline{l_{2}}\vee\ldots\vee\overline{l_{n}}\vee\overline{\mathbf{p}}$.
The proof of Lemma 2 (in Appendix 0.A.3) relies on the duality in Lemma 1. Lemma 2 enables upper-bound construction and paves the way for constructing an instantiation-based Horn upper-bound of a monotonic definition.
@@
The proof of Lemma 3 is in Appendix 0.A.4. Note that the instantiation-based Horn upper-bound of a negative predicate atom $\bar{\mathbf{p}}$ is constructed a monotonic definition of the positive predicate atom $\Sigma_{\mathbf{p}}^{+}$, and vice-versa.
@@
The proof of Theorem 4.3 is in Appendix 0.A.5. Theorem 4.3 states that a lemma-specific Horn upper-bound for a theory lemma $\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}}$ can be constructed by instantiating the monotonic definition using a “witness” assignment $\mathcal{M}_{X}$  [^3] Instead of instantiating a complete assignment on every auxiliary variable in $X$, a partial instantiation is sufficient so long as it determines the assignments on the other variables. . The witness could be obtained by performing SAT solving on the formula $\Sigma_{\mathbf{p}}^{+}|_{\mathcal{M}_{A}^{+}\cup\overline{\mathbf{p}}}$, (where $\mathcal{M}_{A}^{+}$ is the extension of $\mathcal{M}_{A}$ by assigning unassigned input variables in $A$ to $\top$ (Sec. 2.2)). However, in practice, a better approach is to modify the SMMT solver to produce the witness during the derivation of theory lemmas. In Section 5, we provide examples of witnesses for commonly used monotonic predicates.
@@
Continuing the example, let a theory lemma $L:=\mathbf{c}\vee\mathbf{d}\vee\overline{\mathbf{reach_{s}^{t}}}$ be given. To derive a lemma-specific Horn upper-bound for $\Sigma_{\overline{\mathbf{reach_{s}^{t}}}}$, we first obtain a witness $\mathcal{M}_{X}$ by finding a satisfying assignment to the formula $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}\wedge M\wedge\overline{\mathbf{reach_{s}^{t}}}$, where $M:=\{\mathbf{a},\mathbf{b},\overline{\mathbf{c}},\overline{\mathbf{d}},\mathbf{e},\mathbf{f},\mathbf{g},\mathbf{h}\}$ (by assigning the unassigned input variables in $L$ to $\top$). The witness $\mathcal{M}_{X}$ is $\{\mathbf{s},\mathbf{v1},\mathbf{v2},\overline{\mathbf{v3}},\overline{\mathbf{v4}},\overline{\mathbf{t}}\}$. Following the construction in Theorem 4.3, the formula $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}|_{\>\overline{\mathbf{reach_{s}^{t}}}\cup\mathcal{M}_{X}}$ simplifies to two (unit) clauses: $\overline{\mathbf{c}}$ and $\overline{\mathbf{d}}$ (from clauses (2) and (6) in Ex. 1), which can be visualized as the cut in Fig. 2 (right). The lemma-specific Horn upper bound $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}|_{\>\overline{\mathbf{reach_{s}^{t}}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{reach_{s}^{t}}}$ is, therefore, $\overline{\mathbf{c}}\wedge\overline{\mathbf{d}}\Rightarrow\overline{\mathbf{reach_{s}^{t}}}$, which in this example is already CNF, but more generally, we would introduce two literals to encode the implication: $\{\mathbf{c}\vee\overline{\mathbf{l_{1}}},\;\mathbf{d}\vee\overline{\mathbf{l_{2}}},\;\mathbf{l_{1}}\vee\mathbf{l_{2}}\vee\overline{\mathbf{reach_{s}^{t}}}\}$. The lemma-specific Horn upper-bound is dual-Horn and implies the theory lemma $L$ by unit propagation.
@@
In summary, to efficiently verify SMMT theory lemmas, we propose the following approach: (1) define the propositional definitions (in CNF) for the atoms of theory predicates; (2) transform the definitions into monotonic definitions offline; (3) during proof checking, approximate a proof-specific Horn definition (if not already Horn) from the constructed monotonic definition using theory lemmas in the proof; (4) combine the proof-specific definition together and verify the proof via RUP. The only theory-specific, trusted foundation for the proof is the definition for the theory atoms. Appendix. 0.B visualizes this workflow.
@@
The only trusted basis of our approach are the propositional definitions of theory atoms. For the monotonic theories we studied, as discussed in the section 5, the definitions are intuitively understandable and therefore sufficiently trustworthy. For other monotonic theories, the propositional definitions can be validated using techniques from hardware validation and verification, such as simulation, model checking, and equivalence checking.
@@ 5 Example Propositional Definitions
In this section, we illustrate the monotonic definitions for the most commonly used monotonic predicates. Due to space constraints, we present only graph reachability here in detail, and only sketch bit-vector comparison and summation, and max-flow. Full definitions for those theories are in Appendix 0.C.
@@
Instead of defining the monotonic definition for the negative predicate atom $\overline{\mathbf{reach^{v}_{u}}}$, we construct its proof-specific definition from the monotonic definition of the positive predicate atom $\mathbf{reach^{v}_{u}}$. For each theory lemma in the proof, the witness for constructing the lemma-specific Horn upper-bound is the reachability status ($reach^{i}$) of every vertex $i\in V$, which can efficiently computed by traversing the graph in linear time.
@@
Bit-Vector Summation and Comparison (sketch): These are just Tseitin encodings of ripple-carry adders, combined with the comparison theory above. We obtain the definitions for positive and negative theory atoms using the monotonic transformation defined in Definition 6. The resulting positive and negative definitions introduce two sets of auxiliary bit-vectors, $\vec{A^{\prime}}$ and $\vec{B^{\prime}}$, that mirror the two input bit-vectors sets $\vec{A}$ and $\vec{B}$, respectively. Roughly speaking (please see Appendix 0.C.2 for the precise definitions), the definitions captures the following relationships:
$\forall i\in[1,|\vec{A}|].\vec{A}[i]\geq\vec{A^{\prime}}[i]\wedge\forall j\in[1,|\vec{B}|].\vec{B}[j]\leq\vec{B^{\prime}}[j]\wedge\sum\vec{A^{\prime}}\geq\sum\vec{B^{\prime}}\Rightarrow\sum\vec{A}\geq\sum\vec{B}$
$\forall i\in[1,|\vec{A}|].\vec{A}[i]\geq\vec{A^{\prime}}[i]\wedge\forall j\in[1,|\vec{B}|].\vec{B}[j]\leq\vec{B^{\prime}}[j]\wedge\sum\vec{A^{\prime}}<\sum\vec{B^{\prime}}\Rightarrow\sum\vec{A}<\sum\vec{B}$
Note that the definitions are not Horn, and they require witnesses on $\vec{A^{\prime}}$ and $\vec{B^{\prime}}$ to construct lemma-specific Horn definitions. The witnesses are computed by SMMT solvers.

### 04 Apr 2024 v2

@@ Abstract
{ProceedingsVersion}
An extended version of this paper, which includes appendices with proofs and additional results, is available at https://doi.org/10.48550/arXiv.2401.10703 {ExtendedVersion} This is an extended version of a paper published in TACAS 2024. It is essentially identical to that paper, except that it includes appendices with proofs and additional results. However, it does not incorporate changes (if any) made by the publisher, and is not the “Version of Record”. The Version of Record is available at http://dx/doi.org/TBD. Please cite that paper instead of this one, unless you are referencing specific details in the appendices.
@@ 1 Introduction
  • DRAT is a desirable proof format. (Here, we include related formats like DRUP [27], GRIT [19], and LRAT [18]. DRAT is explained in Sec. 2.1.) For an independent assurance of correctness, the proof checker is the critical, trusted component, and hence must be as trustworthy as possible. For (propositional) SAT, the community has coalesced around the DRAT proof format [37], for which there exist independent, efficient proof checkers [37], mechanically verified proof checkers [38], and even combinations that are fast as well as mechanically proven [18]. The ability to emit DRAT proof certificates has been required for solvers in the annual SAT Competition since 2014.
@@
We implemented our method in the MonoSAT SMMT solver [10]. For evaluation, we use two sets of benchmarks derived from practical, industrial problems: multilayer escape routing [11], and cloud network reachability [2].111 Available at https://github.com/NickF0211/MonoProof. Our results show minimal runtime overhead on the solver (geometric mean slowdown 7.4%, worst-case 28.8% in our experiments), and we generate and check proofs for many problem instances that are otherwise intractable.
@@ 2.1 Propositional SAT and DRAT
DRAT [37] is a much more compact and efficient system for proving unsatisfiability. It is based on reverse unit propagation (RUP), which we explain here.222 RUP is all we use in this paper. RAT is a superset of RUP, by essentially doing one step of resolution as a “lookahead” before checking RUP of the resolvents. The “D” in DRAT stands for “deletion”, meaning the proof format also records clause deletions. A unit clause is a clause containing one literal. If $L$ is the set of literals appearing in the unit clauses of a formula $F$, the unit clause rule computes $F|_{L}$, and the repeated application of the unit clause rule until a fixpoint is called unit propagation (aka Boolean constraint propagation). Given a clause $C$, its negation $\overline{C}$ is a set of unit clauses, and we denote by $F\vdash_{1}C$ if $F\wedge\overline{C}$ derives a conflict through unit propagation. Notice that $F\vdash_{1}C$ implies $F\models C$, but is computationally easy to check. The key insight [24] behind RUP is that modern CDCL SAT solvers make progress by deriving learned clauses, whose redundancy is, by construction, checkable via unit propagation. Proof generation, therefore, is essentially just logging the sequence of learned clauses leading to $\bot$, and proof checking is efficiently checking $\vdash_{1}$ of the relevant learned clauses.
@@ 2.2 SAT Modular Monotonic Theories (SMMT)
Given a monotonic predicate $p$ over input $A$, we will use boldface $\mathbf{p}$ as the predicate atom for $p$, i.e., the predicate atom is a Boolean variable in the CNF encoding of the theory, indicating whether $p(A)$ is true or not. The theory of $\mathbf{p}$ is the set of valid implications in the form of $\mathcal{M}_{A}\Rightarrow\mathbf{p}$ where $\mathcal{M}_{A}$ is a partial assignment over $A$.
@@ 4.2 Monotonic Definitions
{ProceedingsVersion}
The proof of Theorem 4.2 is in the extended version of this paper. {ExtendedVersion} The proof of Theorem 4.2 is in Appendix 0.A.1. The correctness relies on the fact that the predicate $p$ is indeed monotonic, and that our propositional definitions need only be one-sided. If the monotonic definition is already in Horn theory, it can be used directly verify theory lemmas via RUP; otherwise, we proceed to Horn approximation, described next.
@@ 4.3 Instantiation-Based, Proof-Specific Horn Definition
Let $p$ be a monotonic predicate over Boolean arguments $A$. Suppose $\Sigma_{\mathbf{p}}$ and $\Sigma_{\bar{\mathbf{p}}}$ are positive and negative propositional definitions, respectively. For every assignment $\mathcal{M}$ to the variables in $A$:
@@
{ProceedingsVersion}
The proof of Lemma 1 is in the extended version of this paper. {ExtendedVersion} The proof of Lemma 1 is presented in Appendix 0.A.2. The duality of the positive ($\Sigma_{\mathbf{p}}$) and negative ($\Sigma_{\bar{\mathbf{p}}}$) definitions allows us to over-approximate positive (negative) definitions by instantiating the negative (positive) definitions.
@@
Let a predicate $p$ over input $A$ and a positive definition $\Sigma_{\mathbf{p}}$ be given. For any partial assignment $M^{\prime}$ over $var(\Sigma_{\mathbf{p}})\setminus(var(p)\cup A)$,   $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\Rightarrow\overline{\mathbf{p}}$ is an over-approximation of $\Sigma_{\overline{\mathbf{p}}}$.    333 Note that $\Sigma_{\mathbf{p}}|_{M^{\prime}}$ is encoded in CNF, so to compactly (i.e., linear-size) encode $\Sigma_{\mathbf{p}}|_{M^{\prime}}\Rightarrow\overline{\mathbf{p}}$ in CNF, we introduce a new literal $l_{i}$ for each clause $C_{i}\in\Sigma_{\mathbf{p}}|_{M^{\prime}}$, create clauses $\overline{c_{ij}}\vee\overline{l_{i}}$ for each literal $c_{ij}\in C_{i}$, and add clause $l_{1}\vee l_{2}\vee\ldots\vee l_{n}\vee\overline{\mathbf{p}}$.
{ProceedingsVersion}
The proof of Lemma 2 (in the extended paper) relies on the duality in Lemma 1. {ExtendedVersion} The proof of Lemma 2 (in Appendix 0.A.3) relies on the duality in Lemma 1. Lemma 2 enables upper-bound construction and paves the way for constructing an instantiation-based Horn upper-bound of a monotonic definition.
@@
{ProceedingsVersion}
(Proof in the extended paper.) {ExtendedVersion} The proof of Lemma 3 is in Appendix 0.A.4. Note that the instantiation-based Horn upper-bound of a negative predicate atom $\bar{\mathbf{p}}$ is constructed from a monotonic definition of the positive predicate atom $\Sigma_{\mathbf{p}}^{+}$, and vice-versa.
@@
{ProceedingsVersion}
(Proof in the extended paper.) {ExtendedVersion} The proof of Theorem 4.3 is in Appendix 0.A.5. Theorem 4.3 states that a lemma-specific Horn upper-bound for a theory lemma $\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}}$ can be constructed by instantiating the monotonic definition using a “witness” assignment $\mathcal{M}_{X}$.  [^4] Instead of instantiating a complete assignment on every auxiliary variable in $X$, a partial instantiation is sufficient so long as it determines the assignments on the other variables. The witness could be obtained by performing SAT solving on the formula $\Sigma_{\mathbf{p}}^{+}|_{\mathcal{M}_{A}^{+}\cup\overline{\mathbf{p}}}$, (where $\mathcal{M}_{A}^{+}$ is the extension of $\mathcal{M}_{A}$ by assigning unassigned input variables in $A$ to $\top$ (Sec. 2.2)). However, in practice, a better approach is to modify the SMMT solver to produce the witness during the derivation of theory lemmas. In Section 5, we provide examples of witnesses for commonly used monotonic predicates.
@@
Continuing the example, let a theory lemma $L:=\mathbf{c}\vee\mathbf{d}\vee\overline{\mathbf{reach_{s}^{t}}}$ be given. To derive a lemma-specific Horn upper-bound for $\Sigma_{\overline{\mathbf{reach_{s}^{t}}}}$, we first obtain a witness $\mathcal{M}_{X}$ by finding a satisfying assignment to the formula $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}\wedge M\wedge\overline{\mathbf{reach_{s}^{t}}}$, where $M:=\{\mathbf{a},\mathbf{b},\overline{\mathbf{c}},\overline{\mathbf{d}},\mathbf{e},\mathbf{f},\mathbf{g},\mathbf{h}\}$ (by assigning the unassigned input variables in $L$ to $\top$). Since $M$ is a complete assignment to the edge variables, the graph is fully specified, and a suitable witness $\mathcal{M}_{X}$ can be efficiently computed using a standard graph-reachability algorithm, to compute the reachability status of each vertex. The witness $\mathcal{M}_{X}$ is $\{\mathbf{s},\mathbf{v1},\mathbf{v2},\overline{\mathbf{v3}},\overline{\mathbf{v4}},\overline{\mathbf{t}}\}$. Following the construction in Theorem 4.3, the formula $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}|_{\>\overline{\mathbf{reach_{s}^{t}}}\cup\mathcal{M}_{X}}$ simplifies to two (unit) clauses: $\overline{\mathbf{c}}$ and $\overline{\mathbf{d}}$ (from clauses (2) and (6) in Ex. 1), which can be visualized as the cut in Fig. 2 (right). The lemma-specific Horn upper bound $\Sigma_{\mathbf{reach_{s}^{t}}}^{h}|_{\>\overline{\mathbf{reach_{s}^{t}}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{reach_{s}^{t}}}$ is, therefore, $\overline{\mathbf{c}}\wedge\overline{\mathbf{d}}\Rightarrow\overline{\mathbf{reach_{s}^{t}}}$, which in this example is already CNF, but more generally, we would introduce two literals to encode the implication: $\{\mathbf{c}\vee\overline{\mathbf{l_{1}}},\;\mathbf{d}\vee\overline{\mathbf{l_{2}}},\;\mathbf{l_{1}}\vee\mathbf{l_{2}}\vee\overline{\mathbf{reach_{s}^{t}}}\}$. The lemma-specific Horn upper-bound is dual Horn and implies the theory lemma $L$ by unit propagation.
@@
In summary, to efficiently verify SMMT theory lemmas, we propose the following approach: (1) define the propositional definitions (in CNF) for the atoms of theory predicates; (2) transform the definitions into monotonic definitions offline; (3) during proof checking, approximate a proof-specific Horn definition (if not already Horn) from the constructed monotonic definition using theory lemmas in the proof; (4) combine the proof-specific definition together and verify the proof via RUP. The only theory-specific, trusted foundation for the proof is the definition for the theory atoms. {ProceedingsVersion} (The extended version of this paper contains a figure to help visualize this workflow.) {ExtendedVersion} Appendix 0.B visualizes this workflow.
@@
The only trusted basis of our approach are the propositional definitions of theory atoms. For the monotonic theories in the section 5, we considered the definitions intuitively understandable, and therefore sufficiently trustworthy. But to further increase confidence, propositional definitions can be validated using techniques from hardware validation/verification, e.g., simulation to sanity-check general behavior, equivalence checking against known-good circuits, etc.
@@ 5 Example Propositional Definitions
In this section, we illustrate the monotonic definitions for the most commonly used monotonic predicates. Due to space constraints, we present only graph reachability here in detail, and only sketch bit-vector comparison and summation, and max-flow. {ProceedingsVersion} Full definitions for those theories are in the extended version of this paper. {ExtendedVersion} Full definitions for those theories are in Appendix 0.C.
@@
Instead of defining the monotonic definition for the negative predicate atom $\overline{\mathbf{reach^{v}_{u}}}$, we construct its proof-specific definition from the monotonic definition of the positive predicate atom $\mathbf{reach^{v}_{u}}$. For each theory lemma in the proof, the witness for constructing the lemma-specific Horn upper-bound is the reachability status ($reach^{i}$) of every vertex $i\in V$, which is efficiently computed in the SMMT solver using standard graph-reachability algorithms.
@@
Bit-Vector Summation and Comparison (sketch): These are basically Tseitin encodings of ripple-carry adders, combined with the comparison theory above — using Def. 6 to handle the fact that the the Tseitin encodings of the XOR gates in the adders are non-monotonic with respect to the input bit-vectors. The resulting propositional definitions are not Horn, so we use witnesses to construct lemma-specific Horn definitions. The witnesses come from the SMMT solver maintaining lower and upper bounds on the possible values of the bit-vectors, e.g., a witness for $\sum{\vec{A}}\geq\sum{\vec{B}}$ are lower bounds for the vectors in $\vec{A}$ and upper bounds for the vectors in $\vec{B}$ such that their sums make the inequality true. (Mutadis mutandis for the negative witness.)